# Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning

### v1

@@ A.1 Experiments Setup
A.1.1 Datasets
We conduct experiments on the FewNERD dataset Ding et al. (2021) which is annotated with 8 coarse-grained and 66 fine-grained entity types. We evaluate our method on the two settings of FewNERD: 1) FewNERD-INTRA, which pertains to a scenario wherein entities within the training set (source domain), validation set, and test set (target domain) are exclusively associated with distinct coarse-grained types. 2) FewNERD-INTER, where only fine-grained entity types that do not intersect in different sets. Furthermore, we follow the settings of ESD Wang et al. (2022a) and adopt $N$-way $K\sim 2K$-shot sampling method to construct tasks.
A.1.2 Parameter Settings
We use grid search for hyperparameter settings and we train our model for 1,000 steps and choose the best model on validation dataset for testing. We use the pretrained language model uncased BERT-base as our encoder. In the entity span detection module, we adopt the BIOES tags. We use the AdamW optimizer with a learning rate of 3e-5 and 0.01 linear warmup steps. The $K$ in KNN is set to 10 and the max-loss coefficient $\lambda$ of KNN is 0.1 when adding KNN. The batch size is set to 32, the max sequence length is set to 128 and we use a dropout probability of 0.2.
Models FewNERD-INTRA FewNERD-INTER
1-2 shot 5-10 shot 1-2 shot 5-10 shot
5 way 10 way 5 way 10 way 5 way 10 way 5 way 10 way
MsFNER 0.3442 0.3808 0.5616 1.0546 0.3482 0.3806 0.6708 1.2266
ChatGPT 4.5581 5.4897 8.1842 9.3481 3.2633 8.0579 4.7755 5.3409
Table 2: The comparison of inference time between MsFNER and ChatGPT.
A.1.3 Baselines
@@
  • ProtoBERT Fritzler et al. (2019) employs the prediction of query labels by leveraging the similarity between the BERT hidden states derived from the support set and the query tokens.
  • CONTAINER Das et al. (2022) employs token-level contrastive learning to train BERT as the token embedding function. Subsequently, BERT undergoes finetuning on the support set, followed by the application of a nearest neighbor method during the inference phase.
  • NNShot Yang and Katiyar (2020) bears resemblance to ProtoBERT in its methodology, as it conducts predictions utilizing nearest neighbor algorithms.
  • StructShot Yang and Katiyar (2020) incorporates an augmented Viterbi decoder into the inference process atop NNShot’s framework.
  • ESD Wang et al. (2022a) formulates few-shot sequence labeling as a span-level matching problem and decomposes it into span-related procedures.
  • MAML-ProtoNet Ma et al. (2022) proposes a decomposed meta-learning approach for few-shot NER, which tackles few-shot span detection and entity typing using MAML to facilitate rapid adaptation to novel entity classes and proposes MAML-ProtoNet for improved representation of entity spans.
  • ChatGPT chatgpt (2023) is an advanced neural language model developed by OpenAI, which is trained on diverse corpora of text and could generate human-like text.
A.2 ChatGPT Prompt Template
@@
In the task description, we initially elucidate the few-shot Named Entity Recognition (NER) task required from ChatGPT. Subsequently, we provide ChatGPT with explicit instructions for implementing this task, including delineation of the entity types involved. To facilitate easier interpretation of model output, we utilize a structured output JSON format, a domain in which ChatGPT exhibits proficiency. The template is as follows:
Your task is to perform fewshot Named Entity Recognition. You could perform the task by the following actions:
@@
    Output a JSON object that contains the following keys: text, entity_list. Note that each item in entity_list is a dictionary with "entity" and "type" as keys.
@@
In the template above, the \$ent_list denotes the list of entity types corresponding to $N$-way. The \$case_input comprises the sentences in the support set corresponding to $K$-shot, while the \$case_output represents the labels in the support set. Since ChatGPT is better at structured output, we use JSON format for output, with each output format being: {"entity": entity text ; "type": entity type}. The output will be merged into a list. The \$input_text denotes the query data. When we input this prompt to ChatGPT, it will return a list of relevant entity-type dictionaries.
A.3 Model Efficiency
To evaluate the efficiency of MsFNER, we compute the average inference time of MsFNER and ChatGPT in different settings. And the results are shown in Table 2, where we can see that MsFNER is on the millisecond level while the ChatGPT is on the second level. The time consumption is related to the amount of input data. In addition, the time consumption of ChatGPT will increase due to network impact.
A.4 training process
As the Figure 3 shows, the training loss on training dataset is smoothly declining within the same source types. But when finetuning on the validation set, there are some fluctuations due to the cross domain transfer of knowledge. And with the training time increasing, the overfitting is occurred slightly on validation dataset because of few supervised samples.
Figure 3: The performance of training process on FewNERD 5-way 5~10-shot. (a) is the evaluation F1 score when finetuning on the validation dataset. (b) is the training loss on train dataset.

### v2

@@ III-A Experiments Setup
III-A1 Datasets
We conduct experiments on the FewNERD dataset [20] which is annotated with 8 coarse-grained and 66 fine-grained entity types. We evaluate our method on the two settings of FewNERD-INTRA and FewNERD-INTER, which are with the different grained types. Furthermore, we follow the settings of ESD [10] and adopt $N$-way $K\sim 2K$-shot sampling method to construct tasks. Another dataset we use for testing our method is FewAPTER, which is built by us based on the cybersecurity threat intelligence dataset APTER[21].
Specifically, APTER is proposed by Wang et al.[21] within general data distributions, and it predefines 37 entity types under domain expert knowledge. In this paper, we refine and enhance the original APTER dataset, consolidating its entity types into 21 categories, resulting in a total of 28,250 entities. Inspired by the design methodology of the FewNERD dataset, we follow the knowledge transfer principle (from original entity type data $\rightarrow$ to new entity type data) and construct the FewAPTER dataset based on APTER for few-shot entity recognition research in the cybersecurity domain. Considering the data volume of each entity type, the 21 entity types are categorized into three subsets: the training set with 7 types {VULID, FILE, SECTEAM, LOC, TOOL, APT, ACT}, the validation set with 7 types {SHA1, OS, URL, MAL, DOM, PROT, TIME}, and the test set with 7 types {ENCR, VULNAME, IP, MD5, SHA2, IDTY, EMAIL}.
While constructing the dataset, it is necessary that the entity types in the training, validation, and test sets do not overlap. Therefore, a maximum of 7-way can be set. However, due to the imbalanced data distribution and the possibility of multiple entity types occurring in a single sentence, setting the dataset to 7-way limits the diversity of knowledge learned, making it challenging to adapt to and generalize for new tasks. To address this, this paper constructs four $N$-way $K$-shot datasets: {4-way 1-shot, 4-way 3-shot, 6-way 1-shot, 6-way 3-shot}. Each $N$-way $K$-shot dataset includes 4000 training samples, along with 800 validation samples and 800 test samples. During the construction of each $N$-way $K$-shot dataset (for training, validation, and testing), each task, denoted as $S^{(i)}$ or $Q^{(i)}$, is randomly selected from the corresponding dataset, consisting of $N$ entity types and $K$ examples. For example, when constructing a task $S^{(i)}_{source}$ or $Q^{(i)}_{source}$ for the 6-way 3-shot training set, 6 entity types are randomly chosen from the 7 available types, with 3 entity examples selected for each entity type.
III-A2 Parameter Settings
We use grid search for hyperparameter settings and we train our model for 1,000 steps and choose the best model on validation dataset for testing. We use the pretrained language model uncased BERT-base as our encoder. In the entity span detection module, we adopt the BIOES tags. We use the AdamW optimizer with a learning rate of 3e-5 and 0.01 linear warmup steps. The $K$ in KNN is set to 10 and the coefficient $\lambda$ of KNN is 0.1 when adding KNN. The batch size is set to 32, the max sequence length is set to 128 and we use a dropout probability of 0.2.
III-B Baselines
@@
  • ProtoBERT [5] employs the prediction of query labels by leveraging the similarity between the BERT hidden states derived from the support set and the query tokens.
  • CONTAINER [9] employs BERT as the encoder and designs a token-level contrastive method to enhance the token embedding for few-shot NER.
  • NNShot [8] bears resemblance to ProtoBERT in its methodology, as it conducts predictions utilizing nearest neighbor algorithms.
  • StructShot [8] incorporates an augmented Viterbi decoder into the inference process atop NNShot’s framework.
  • ESD [10] formulates few-shot sequence labeling as a span-level matching problem and decomposes it into span-related procedures.
  • MAML-ProtoNet [12] proposes a meta-learning approach with fusing MAML and prototype network for few-shot NER, which facilitates rapid adaptation to novel entity types.
  • BDCP [18] is a few-shot NER method with boundary discrimination and correlation purification, which explores the robustness of the model’s cross domain transfer learning ability in textual adversarial attack scenarios.
  • ChatGPT [22] is an advanced large language model (LLM) developed by OpenAI, which is trained on diverse corpora of text and could generate human-like text.
III-C ChatGPT Prompt Template
@@
In the task description, we initially elucidate the few-shot NER task required from ChatGPT. Subsequently, we provide ChatGPT with explicit instructions for implementing this task, including delineation of the entity types involved. To facilitate easier interpretation of model output, we utilize a structured output JSON format, a domain in which ChatGPT exhibits proficiency. The template is as follows:
Your task is to perform few-shot Named Entity Recognition. You could perform the task by the following actions:
@@
    Output a JSON object that contains the following keys: text, entity_list. Note that each item in entity_list is a dictionary with “entity” and “type” as keys.
@@
In the template above, the \$ent_list denotes the list of entity types corresponding to $N$-way. The \$case_input comprises the sentences in the support set corresponding to $K$-shot, while the \$case_output represents the labels in the support set. Since ChatGPT is better at structured output, we use JSON format for output, with each output format being: {“entity”: entity text ; “type”: entity type}. The output will be merged into a list. The \$input_text denotes the query data. When we input this prompt to ChatGPT, it will return a list of relevant entity-type dictionaries.
III-D Main results
Table I illustrates main results of different methods on FewNERD dataset and Table III shows results on the FewAPTER dataset. These results demonstrate that MsFNER significantly outperforms previous few-shot NER approaches. We have several findings: 1) MsFNER attains average F1 improvements of 2.65% and 4.44% on FewNERD-INTRA and FewNERD-INTER, respectively, compared to the previous best few-shot NER method, MAML-ProtoNet. Furthermore, the greater improvement is on the FewAPTER dataset, which achieves an average F1 score improvement of 11.42%. 2) Although MsFNER outperforms the ChatGPT on the FewNERD dataset, it is slightly inferior to the LLM on FewAPTER. This is due to the characteristics of the FewAPTER dataset, which enable the model to exhibit strong learning capabilities within a specific domain, but with slightly reduced generalization ability for cross-domain learning. 3) Regardless of whether it is ChatGPT or small models, increasing the number of $K$-shot yields superior performance, while an increase in the number of $N$-way results in worse performance. This aligns with general human intuition: given a fixed number of entity types, having more examples leads to better learning accuracy; conversely, with a fixed number of examples, increasing the number of entity types makes learning more challenging. 4) From the performance evaluation of these results, it can be seen that the performance of LLM is quite stable, which significantly indicates that its influence from data and domain changes is relatively limited. This is primarily attributed to the large number of parameters in the model, which, unlike few-shot learning models, exhibits lower sensitivity to the number of training samples.
III-E Ablation Study
To investigate the contributions of different modules of MsFNER, we conduct the ablation study by removing each of them at a time. The results are shown at the bottom of Table I and Table III. Firstly, we remove the contrastive learning module. In the 1-shot experiment of this setting, KNN is unnecessary as there is only one sample in the support dataset. We can see that the removal of contrastive learning results in an average decrease of 0.905% on FewNERD and 0.745% on FewAPTER, which indicates the significant impact and necessity of contrastive learning for enhancing representation. Secondly, we remove the KNN module and reserve the contrastive learning module. The removal of KNN drops average 0.524% on FewNERD and 0.635% on FewAPTER. We can see that the impact of contrastive learning and KNN changes with the number of $N$-way and $K$-shot. Removing either of them will lead to a decrease in the performance of MsFNER.
III-F Model Efficiency
To comprehensively evaluate MsFNER and ChatGPT, we compute their average inference time(s) in different settings, and the results are shown in Table II. From the table, we can see that MsFNER is much faster than its LLM peer ChatGPT. According to the Occam’s Razor, we think that MsFNER is a better choice for few-shot NER in application.
TABLE III:  F1 scores with standard deviations(%) on FewAPTER
Models FewAPTER
1 shot 3 shot
4 way 6 way 4 way 6 way
ESD† 33.83±0.57 29.97±0.42 41.60±0.27 39.08±0.72
CONTAINER† 29.82 31.17 51.71 50.65
BDCP† 44.65±0.97 36.55±1.02 51.43±0.63 39.55±1.74
MAML-ProtoNet† 48.41±0.31 39.32±0.19 49.51±0.47 37.16±0.11
ChatGPT‡ 59.14 57.65 61.14 57.68
MsFNER 57.55±0.39 49.90±0.90 61.36±0.59 51.26±1.38
w/o cl* 60.78±0.30 50.35±0.24
w/o KNN 56.47±0.69 49.23±0.80 61.21±0.17 50.62±1.23
III-G Training and Validation Process
[…]
Figure 3: The performance of training and validation stage on two different settings
IV Conclusion
In this paper, we propose MsFNER, a novel approach to addressing the challenges of a large number of negative samples and computational overhead in the domain of few-shot named entity recognition. MsFNER is implemented in three stages: training, fine-tuning, and inference. During the training and fine-tuning stages, we first detect the entity spans and subsequently recognize the entity types. To enhance the learning of entity representations, we introduce a contrastive learning module. In the inference stage, we replace the contrastive learning module with a KNN module to further improve inference performance. The final inference output is determined by combining the KNN results and the type classification predictions. We conduct experiments on the FewNERD and FewATPER datasets to evaluate the performance of MsFNER. Additionally, we compare our approach with ChatGPT-3.5 on few-shot tasks, and the results demonstrate that MsFNER can achieve performance comparable to ChatGPT-3.5 to some extent.